# Label Distribution Shift-Aware Prediction Refinement for Test-Time Adaptation

### v1

@@ 2 Label distribution shifts on TTA
For more detailed explanations and proofs of these confusion patterns, please refer to Appendix D. Through experiments with real datasets, we reaffirm the impact of label distribution shifts on a BN-adapted classifier. We consider a BN-adapted classifier originally trained on a class-balanced CIFAR-10 dataset [25] and later tested using CIFAR-10-LT and CIFAR-10C-LT test datasets, which include label distribution shifts. CIFAR-10C [26] serves as a benchmark for assessing model robustness against 15 different predefined types of corruptions, such as Gaussian noise. To analyze the effects of label distribution shifts, we define the number of samples for class $k$ as $n_{k}=n(1/\rho)^{k/(K-1)}$, where $n$ is the number of samples for the head class and $\rho$ is the class imbalance ratio. In Fig. 3, we observe substantial performance drops in test accuracy for BNAdapt (orange) as the class imbalance ratio $\rho$ increases. The BNAdapt exhibits even worse performance than NoAdapt (blue) when $\rho=100$.
In Fig. 1, we provide more detailed analysis for misclassification patterns by presenting confusion matrices illustrating the impact of two distinct long-tailed distributions (each row) across four types of image corruptions (each column, including clean, Gaussian noise, Defocus blur, and Snow). These matrices show the fraction of test samples from class $i$ classified into class $j$. The trends observed in these real experiments are similar to those analyzed in our toy example. Firstly, significant accuracy degradation is evident, especially for head classes (with smaller/larger class indices in the 1st/2nd row, respectively). Secondly, the confusion patterns remain consistent across different corruption types and reflect class-wise relationships; for example, more frequent confusion occurs between classes with similar characteristics, such as airplane & ship (0 and 8), automobile & truck (1 and 9), and cat & dog (3 and 5). Additionally, we present confusion matrices for CIFAR-10C-LT with various levels of label distribution shifts ($\rho=1,10,100$) in Figures 7-9 of Appendix F, revealing increasingly pronounced class-wise confusion patterns as the imbalance ratio $\rho$ rises from 1 to 100.
@@ D.1 Motivating toy example - confusion patterns reflecting the class relationship by BNAdapt
The effects of test-time label distribution shift can be consistently observed not only in this toy example but also in real dataset experiments, e.g., CIFAR-10C-LT (Section 2).
@@ Appendix E Confusion matrices of BN-adapted classifiers on CIFAR-10C-LT
In Figure 1, we present confusion matrices of the BN-adapted classifiers for three different corruption types (Gaussian noise, defocus blur, snow) from each of the three corruption categories (noise, blur, and weather) among 15 pre-defined corruptions. In Figure 7-9, we present confusion matrices for all 15 corruptions including clean CIFAR-10 test dataset with various levels of label distribution shifts ($\rho=1,10,100$), respectively. While test accuracy may not perfectly be matched across the corruptions, we can observe that (1) the accuracy in head classes (with smaller class index) is significantly decreased and (2) confusion patterns tend to be consistent across different corruptions as we described in Section 2. Additionally, Figure 7-9 reveal increasingly pronounced class-wise confusion patterns as the imbalance ratio $\rho$ rises from 1 to 100. Based on these observations, we conjecture that the prediction refinement module trained only on the training data at the intermediate time remains consistently effective in modifying predictions for the test datasets.
@@ Appendix F Prediction refinement module outputs on CIFAR-10C-LT
In Figures 10-12, we present the $W$ and $b$, generated by $g_{\phi}$ when inputting the averaged pseudo label distribution and prediction deviation computed on the whole test data for CIFAR-10C-LT under Gaussian noise with different $\alpha$ values (0, 0.1, and 1) in equation 3. The (a), (b), and (c) of each figure show the outputs of $g_{\phi}$ when class imbalance $\rho$ is set to 1, 10, and 100, respectively. In Figures 10-12, we commonly observe that both $W$ and $b$ assign high values for the head classes (classes 0-1). This prediction refinement of DART can correct the predictions of instances from the head classes that are misclassified to other classes, recovering the reduced test accuracy of the head classes (shown in Figure 1). As shown in Figure 13 of Appendix G, this prediction refinements scheme of DART can modify the averaged pseudo label distribution to be closely matched with the ground truth label distribution, which is originally significantly disagreed.

### v2

@@ 2 Label distribution shifts on TTA
More additional explanations for these confusion patterns including the symbol table, illustrations, and detailed proofs are described in Appendix D. Through experiments with real datasets, we reaffirm the impact of label distribution shifts on a BN-adapted classifier. We consider a BN-adapted classifier originally trained on a class-balanced CIFAR-10 dataset [25] and later tested using CIFAR-10-LT and CIFAR-10C-LT test datasets, which include label distribution shifts. CIFAR-10C [26] serves as a benchmark for assessing model robustness against 15 different predefined types of corruptions, such as Gaussian noise. To analyze the effects of label distribution shifts, we define the number of samples for class $k$ as $n_{k}=n(1/\rho)^{k/(K-1)}$, where $n$ is the number of samples for the head class and $\rho$ is the class imbalance ratio. In Fig. 3, we observe substantial performance drops in test accuracy for BNAdapt (orange) as the class imbalance ratio $\rho$ increases. The BNAdapt exhibits even worse performance than NoAdapt (blue) when $\rho=100$.
In Fig. 1, we provide more detailed analysis for misclassification patterns by presenting confusion matrices illustrating the impact of two distinct long-tailed distributions (each row) across four types of image corruptions (each column, including clean, Gaussian noise, Defocus blur, and Snow). These matrices show the fraction of test samples from class $i$ classified into class $j$. The trends observed in these real experiments are similar to those analyzed in our toy example. Firstly, significant accuracy degradation is evident, especially for head classes (with smaller/larger class indices in the 1st/2nd row, respectively). Secondly, the confusion patterns remain consistent across different corruption types and reflect class-wise relationships; for example, more frequent confusion occurs between classes with similar characteristics, such as airplane & ship (0 and 8), automobile & truck (1 and 9), and cat & dog (3 and 5). Additionally, we present confusion matrices for CIFAR-10C-LT with various levels of label distribution shifts ($\rho=1,10,100$) in Figures 9-11 of Appendix F, revealing increasingly pronounced class-wise confusion patterns as the imbalance ratio $\rho$ rises from 1 to 100.
@@ 6 Discussion and Conclusion
6.1 Discussion
Performance on large-scale benchmarks
While DART consistently enhances BN-adapted classifiers on large-scale datasets such as OfficeHome and PACS under test-time label distribution shifts (Table 1), the performance gains by DART may appear less noticeable on large-scale benchmarks compared to small-scale benchmarks. In these settings, the batch size is typically smaller than the total number of classes, so even if the test dataset’s class distribution differs significantly from that of the training dataset, each batch experiences only minimal distribution shifts. Consequently, while DART’s improvement may seem less significant compared to those on small-scale datasets like CIFAR-10C, it still effectively adjusts inaccurate predictions by label distribution shifts. Additionally, as the number of classes increases, the variety of potential class distribution shifts grows dramatically. Therefore, training the prediction refinement module at the intermediate time becomes more challenging on large-scale benchmarks with many classes, such as ImageNet-C (Table 2). This highlights a need for further research on more efficient methods to train the prediction refinement module, especially in extremely large label spaces.
The use of training data during the intermediate time
Although DART employs labeled training data to train its prediction refinement module before the test time, it does not violate TTA protocols and is commonly adopted by various TTA methods. In particular, DART’s intermediate-time training with labeled data does not conflict with TTA guidelines because the training data is exclusively used to refine the prediction module during the intermediate time and is no longer accessible during the test time. Moreover, as described in Section 5, several recent TTA methods [27, 28, 29] also utilize training data before test time to prepare for unknown distribution shifts. Nevertheless, in scenarios where the labeled training data for a given trained model is unclear, obtaining the training data can be challenging. To increase the scalability of DART in these situations, exploring approaches that leverage an auxiliary dataset instead of the original training data may be a promising direction for future work.
6.2 Conclusion
@@
Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00408003 and No. 2021R1C1C11008539).
@@ A.3 Details about intermediate time training
Figure 7: Class distributions of five batches sampled from CIFAR-10 using Dirichlet sampling. These sampled batches demonstrate diverse and varied class distributions, highlighting the effectiveness of Dirichlet sampling in generating a wide range of label distribution shifts.
@@ D.1 Motivating toy example - confusion patterns reflecting the class relationship by BNAdapt
Table 9: Notation and definitions used in Section D.1
Symbol Meaning
$\mathcal{N}(\mu,\sigma^{2}I_{2})$ Multivariate Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$
$\mu_{i}$ Mean vector of class $i$ ($i=1,2,3,4$)
$\sigma^{2}$ Variance shared across all classes
$d,\beta$ Constants controlling class separation ($d>1,\beta>1$)
$\Delta$ Test-time distribution shift vector
$h(\cdot)$ Bayes classifier for prediction
$\bar{h}(\cdot)$ Mean-centered Bayes classifier for prediction
$p_{\text{tr}}(y)$ Class prior probability during training (uniform: $1/4$)
$p_{\text{te}}(y)$ Class prior probability during testing (imbalanced distribution)
$\mu^{\prime}_{i}$ Shifted mean of class $i$ after test-time mean centering
$\Phi(\cdot)$ Cumulative distribution function (CDF) of the standard normal distribution
@@
Figure 8: Illustration of the 4-class Gaussian mixture model from Section D.1. The two lines indicate the Bayes classifier decision boundaries, partitioning the feature space into regions corresponding to different classes. The patterned areas highlight regions where misclassifications occur. Specifically, the red patterned area denotes regions where samples from Class 1 are misclassified as Class 2, while the green patterned area represents regions where samples from Class 1 are misclassified as Class 3. Two notable confusion patterns can be observed when both input data and label distribution shifts occur, and mean centering is applied. First, Class 1, which has a larger sample size, is more prone to misclassification into other classes compared to smaller classes. Second, misclassifications are more likely to occur toward close classes in the feature space. For instance, Class 1 is more frequently confused with the closer class (Class 2) than with other classes.
@@
The illustration of the class-wise confusion pattern caused by test-time input data and label distribution shift in a 4-class Gaussian mixture model can be found in Figure 8. The effects of test-time label distribution shift can be consistently observed not only in this toy example but also in real dataset experiments, e.g., CIFAR-10C-LT (Section 2).
@@ Appendix E Confusion matrices of BN-adapted classifiers on CIFAR-10C-LT
In Figure 1, we present confusion matrices of the BN-adapted classifiers for three different corruption types (Gaussian noise, defocus blur, snow) from each of the three corruption categories (noise, blur, and weather) among 15 pre-defined corruptions. In Figure 9-11, we present confusion matrices for all 15 corruptions including clean CIFAR-10 test dataset with various levels of label distribution shifts ($\rho=1,10,100$), respectively. While test accuracy may not perfectly be matched across the corruptions, we can observe that (1) the accuracy in head classes (with smaller class index) is significantly decreased and (2) confusion patterns tend to be consistent across different corruptions as we described in Section 2. Additionally, Figure 9-11 reveal increasingly pronounced class-wise confusion patterns as the imbalance ratio $\rho$ rises from 1 to 100. Based on these observations, we conjecture that the prediction refinement module trained only on the training data at the intermediate time remains consistently effective in modifying predictions for the test datasets.
@@ Appendix F Prediction refinement module outputs on CIFAR-10C-LT
In Figures 12-14, we present the $W$ and $b$, generated by $g_{\phi}$ when inputting the averaged pseudo label distribution and prediction deviation computed on the whole test data for CIFAR-10C-LT under Gaussian noise with different $\alpha$ values (0, 0.1, and 1) in equation 3. The (a), (b), and (c) of each figure show the outputs of $g_{\phi}$ when class imbalance $\rho$ is set to 1, 10, and 100, respectively. In Figures 12-14, we commonly observe that both $W$ and $b$ assign high values for the head classes (classes 0-1). This prediction refinement of DART can correct the predictions of instances from the head classes that are misclassified to other classes, recovering the reduced test accuracy of the head classes (shown in Figure 1). As shown in Figure 15 of Appendix G, this prediction refinements scheme of DART can modify the averaged pseudo label distribution to be closely matched with the ground truth label distribution, which is originally significantly disagreed.
@@
We experimentally demonstrate the effect of regularization (equation 2) and present the results in Table 10. Without regularization ($\alpha=0$), DART-applied TENT shows approximately 2% lower performance under no label distribution shift ($\rho=1$) compared to other $\alpha$ values. It indicates that DART without regularization makes unnecessary adjustments, harming the subsequent test-time adaptation when there is no class distribution shift. On the other hand, the TTA performance remains stable across all $\rho$s when $\alpha$ is set between 0.01 and 1. This result demonstrates the effectiveness of regularization and the robustness of DART to changes in $\alpha$.
Table 10: Accuracy (%) of DART-applied BNAdapt and TENT with different $\alpha$ values on CIFAR-10C-LT.
Method $\alpha$ CIFAR-10C-LT
$\rho=1$ $\rho=10$ $\rho=100$
BNAdapt + DART 0 85.3 85.5 85.8
0.01 85.4 85.1 85.7
0.1 (used) 85.2 84.7 85.1
0.5 85.3 83.9 84.8
1 85.3 83.5 84.7
TENT + DART 0 84.4 86.2 88.3
0.01 86.1 86.9 88.7
0.1 (used) 86.4 86.8 88.2
0.5 86.6 86.5 87.9
1 86.5 86.2 87.8
@@ G.4 Experimental results using different hyperparameters on CIFAR-10C-LT
Table 15: Performance of DART-applied BNAdapt on CIFAR-10C-LT with different intermediate-time training epochs
CIFAR-10C-LT
# of epochs $\rho=1$ $\rho=10$ $\rho=100$
1 82.0 80.6 75.7
5 85.3 84.5 81.7
10 85.4 84.9 83.3
25 85.3 84.8 84.8
50 (used) 85.2 84.7 85.1
We summarize the experimental results on the effect of training epochs for the prediction refinement module $g_{\phi}$ during the intermediate-time training in Table 15. We set the training epochs for intermediate-time training to 50 for CIFAR-10. To assess the convergence of $g_{\phi}$ training, we conduct experiments by reducing the number of training epochs from 1 to 25. Notably, training for only 10 epochs achieves comparable results to 50 epochs, with just a 1.8% gap observed for CIFAR-10C-LT with $\rho=100$. This demonstrates the efficiency of DART’s intermediate-time training, significantly enhancing its scalability for practical use.